# Fine-Tuning Discrete Diffusion Models via Reward Optimization with Applications to DNA and Protein Design

### v1

@@ 1 Introduction
Diffusion models have gained widespread recognition as effective generative models in continuous spaces, such as image and video generation (Song et al., 2020; Ho et al., 2022). Inspired by seminal works (e.g., Austin et al. (2021); Campbell et al. (2022); Sun et al. (2022)), recent studies (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024) have shown that diffusion models are also highly effective in discrete spaces, including natural language and biological sequence generation (DNA, RNA, proteins). Unlike autoregressive models commonly used in language modeling, diffusion models are particularly well-suited for biological sequences, where long-range interactions are crucial for the physical behavior of molecules arising from those sequences (e.g., the 3D folded structure of RNA or proteins).
@@
While discrete diffusion models effectively capture conditional distributions (e.g., the distribution of sequences given a specific backbone structure in an inverse protein folding design problem (Dauparas et al., 2022; Campbell et al., 2024)), in many applications, especially for therapeutic discovery, we often aim to generate sequences that are both natural-like and optimize a downstream performance objective. For instance, in the inverse folding problem, we may prefer stable protein sequences (i.e., sequences that fold back into stable protein conformations (Widatalla et al., 2024)); for mRNA vaccine production we desire 5’ UTRs that drive high translational efficiency (Castillo-Hair and Seelig, 2021); for gene and cell therapies, we desire regulatory DNA elements, such as promoters and enhancers, that drive high gene expression only in specific cell types (Taskiran et al., 2024); and for natural language we optimize to minimize harmfulness (Touvron et al., 2023).
@@
To address these challenges, our work introduces a fine-tuning approach for well-pretrained discrete diffusion models that maximizes downstream reward functions. Specifically, we aim to optimize these reward functions while ensuring that the generated sequences maintain a high probability under the original conditional distribution (e.g., the distribution of sequences that fold into a given backbone structure). To achieve this, we formulate the problem as a reward maximization task, analogous to reinforcement learning (RL), where the objective function integrates both the reward terms and the KL divergence with respect to the pre-trained discrete diffusion model, which ensures that the generated sequences remain close to the pre-trained model, preserving their naturalness after fine-tuning. To solve this RL problem, we propose a novel algorithm, DRAKES, that enables direct backpropagation of rewards through entire trajectories by making the originally non-differentiable trajectories differentiable using the Gumbel-Softmax trick (Jang et al., 2016).
@@
Our main contribution is an RL-based fine-tuning algorithm, Direct Reward bAcKpropagation with gumbEl Softmax trick (DRAKES), that enables reward-maximizing finetuning for discrete diffusion models (Figure 1). We derive a theoretical guarantee that demonstrates its ability to generate natural and high-reward designs, and demonstrate its performance empirically on DNA and protein design tasks. While similar algorithms exist for continuous spaces (Fan et al., 2023; Black et al., 2023; Uehara et al., 2024), our work is the first, to the best of our knowledge, to address these aspects in (continuous-time) discrete diffusion models. This requires addressing unique challenges, as discrete diffusion models are formulated as continuous-time Markov chains (CTMC), which differ from Brownian motion, and the induced trajectories from CTMC are no longer differentiable, unlike in continuous spaces. Our novel theoretical guarantee also establishes a connection with recent advancements in classifier guidance for discrete diffusion models (Nisonoff et al., 2024).
@@ 2 Related Works
Building on the seminal works of Austin et al. (2021); Campbell et al. (2022), recent studies on masked diffusion models (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024) have demonstrated strong performance in natural language generation. Recent advances in masked discrete diffusion models have been successfully applied to biological sequence generation, including DNA and protein sequences (Sarkar et al., 2024; Campbell et al., 2024). Compared to autoregressive models, diffusion models may be particularly well-suited for biological sequences, which typically yield molecules that fold into complex three-dimensional (3D) structures.
In contrast to these works, our study focuses on fine-tuning diffusion models to optimize downstream reward functions. One application of our approach is the fine-tuning of protein inverse folding generative models to optimize stability, as discussed in Widatalla et al. (2024). However, unlike this prior work, we employ discrete diffusion models as the generative model.
@@
  • Guidance: Techniques such as classifier guidance (Song et al., 2020; Dhariwal and Nichol, 2021) and its variants (e.g., Bansal et al. (2023); Chung et al. (2022); Ho et al. (2022)) introduce gradients from proxy models during inference. However, since gradients are not formally well-defined for discrete states in diffusion, a recent study (Nisonoff et al., 2024) proposed a method specifically designed for discrete diffusion models. Alternative approaches directly applicable to discrete diffusion models include sequential Monte Carlo (SMC)-based methods (Wu et al., 2024; Trippe et al., 2022; Dou and Song, 2024; Cardoso et al., 2023; Phillips et al., 2024). While these guidance-based inference techniques have their own advantages, they generally lead to longer inference times compared to fine-tuned models. We compare our methods against these in terms of generation quality in Section 6.
@@
  • Classifier-free fine-tuning (Ho and Salimans, 2022): This approach constructs conditional generative models, applicable in our setting by conditioning on high reward values. Although not originally designed as a fine-tuning method, it can also be adapted for fine-tuning (Zhang et al., 2023) by adding further controls to optimize. However, in the context of continuous diffusion models, compared to RL-based fine-tuning, several works (Uehara et al., 2024) have shown that conditioning on high reward values is suboptimal, because such high-reward samples are rare. We will likewise compare this approach to ours in Section 6. Lastly, when pretrained models are conditional diffusion models (i.e., $p(x|c)$) and the offline dataset size consisting of triplets $(c,x,r(x))$ is limited, it is challenging to achieve success. Indeed, for this reason, most current RL-based fine-tuning papers (e.g., Fan et al. (2023); Black et al. (2023); Clark et al. (2023)) do not empirically compare their algorithms with classifier-free guidance.
Dirichlet diffusion models for discrete spaces.
@@ 3.1 Diffusion Models on Discrete Spaces
In diffusion models, our goal is to model the data distribution $p_{\mathrm{data}}\in\Delta(\mathcal{X})$ using the training data, where $\mathcal{X}$ represents the domain. We focus on the case where $\mathcal{X}=\{1,2,\cdots,N\}$. The fundamental principle is (1) introducing a known forward model that maps the data distribution to a noise distribution, and (2) learning the time reversal that maps the noise distribution back to the data distribution (detailed in Lou et al. (2023); Sahoo et al. (2024); Shi et al. (2024)).
@@
$\textstyle\frac{dj_{t}}{dt}=Q(t)j_{t},\quad p_{0}\sim p_{\mathrm{data}},$
where $Q(t)\in\mathbb{R}^{N\times N}$ is the generator. Generally, $j_{t}$ is designed so that $p_{t}$ approaches a simple limiting distribution at $t=T$. A common approach is to add Mask into $\mathcal{X}$ and gradually mask a sequence so that the limiting distribution becomes completely masked (Shi et al., 2024; Sahoo et al., 2024).
@@
$\textstyle\frac{dj_{T-t}}{dt}=\bar{Q}(T-t)j_{T-t},\quad\bar{Q}_{y,x}(t)=\begin{cases}\textstyle\frac{j_{t}(y)}{j_{t}(x)}\bar{Q}_{x,y}(t)\,(y\neq x)\\ \textstyle-\sum_{y\neq x}\bar{Q}_{y,x}(t)\,(y=x),\end{cases}$
where $Q_{y,x}(t)$ is a $(y,x)$-entry of a generator $Q(t)$. This implies that if we can learn the marginal density ratio $p_{t}(y)/p_{t}(x)$, we can sample from the data distribution at $t=T$ by following the above CTMC controlled by $\bar{Q}(T-t)$. Existing works (e.g., Lou et al. (2023)) demonstrate how to train this ratio from the training data. Especially when we use masked diffusion models (Sahoo et al., 2024; Shi et al., 2024), we get
$\textstyle\bar{Q}_{y,x}(t)=\begin{cases}\gamma\mathbb{E}[x_{0}=y|x_{t}=\mathrm{Mask}]\quad(y\neq\mathrm{Mask},x_{t}=\mathrm{Mask})\textstyle,\\ -\sum_{z\neq\mathrm{Mask}}\gamma\mathbb{E}[x_{0}=z|x_{t}=\mathrm{Mask}]\quad(y=\mathrm{Mask},x_{t}=\mathrm{Mask})\\ 0\quad(x_{t}\neq\mathrm{Mask})\textstyle\end{cases},$ (1)
@@
When dealing with sequences of length $M$, $x=[x^{\langle 1\rangle},\cdots,x^{\langle M\rangle}]$, we simply consider the factorized rate matrix, i.e., $\bar{Q}_{y,x}=\sum_{i}\bar{Q}_{y^{\langle i\rangle},x}$ (Campbell et al., 2022), thereby avoiding exponential blowup.
@@ 3.2 Goal: Generating Natural Samples While Optimizing Reward Functions
In our work, we consider a scenario with a pretrained discrete diffusion model $p^{\mathrm{pre}}(x|c)\in[\mathcal{C}\to\Delta(\mathcal{X})]$ trained on an extensive dataset and a downstream reward function $r:\mathcal{X}\to\mathbb{R}$. The pretrained diffusion model captures the naturalness or validity of samples. For example, in protein design, $p^{\mathrm{pre}}(\cdot|\cdot)$ could be a protein inverse-folding model that generates amino acid sequences that fold back into the given backbone structure (similar to Campbell et al. (2024)), and $r$ could be a function that evaluates stability. Our objective is to fine-tune a generative model to generate natural-like samples (high $\log p^{\mathrm{pre}}(\cdot|\cdot)$) with desirable properties (high $r(\cdot)$).
@@
We introduce a discrete diffusion model parameterized by $\theta$ from $t=0$ to $t=T$:
@@
The parameter $\theta$ from the pretrained model is denoted by $\theta_{\mathrm{pre}}$. The distribution at time $T$ is denoted as $p^{\mathrm{pre}}(\cdot)$, which approximates the training data distribution $p^{\mathrm{data}}$. We denote an element of the generated trajectory from $t=0$ to $t=T$ by $x_{0:T}$. For simplicity, we assume the initial distribution is a Dirac delta distribution (completely masked state), and we often treat the original pretrained diffusion model as an unconditional model for a single token for notational convenience. In this paper, all of the proofs are in Appendix B.
@@ 4.1 Key Formulation
Perhaps the most obvious starting point for fine-tuning diffusion models to maximize a reward function $r(x_{T})$ is to simply maximize the expected value of the reward under the model’s distribution, i.e., $\mathbb{E}_{x_{0}\sim P^{\theta}}\left[r(x_{T})\right]$, where the expectation is taken over the distribution $P^{\theta}(x_{0:T})$ induced by (2) (i.e., the generator $Q^{\theta}$). However, using only this objective could lead to over-optimization, where the model produces unrealistic or unnatural samples that technically achieve a high reward, but are impossible to generate in reality. Such samples typically exploit flaws in the reward function, for example, by being outside the training distribution of a learned reward or violating the physical assumptions of a hand-engineered physics-based reward (Levine et al., 2020; Clark et al., 2023; Uehara et al., 2024). We address this challenge by constraining the optimized model to remain close to a pretrained diffusion model, which captures the distribution over natural or realistic samples. More specifically, we introduce a penalization term by incorporating the KL divergence between the fine-tuned model $P^{\theta}(x_{0:T})$ and the pretrained diffusion model $P^{\theta_{\mathrm{pre}}}(x_{0:T})$ in CTMC.
@@
$\exp(r(\cdot)/\alpha)p^{\mathrm{pre}}(\cdot).$
@@
A similar theorem has been derived for continuous diffusion models (Uehara et al., 2024, Theorem 1). However, our formulation (3) differs significantly as our framework is based on a CTMC, whereas those works are centered around the Brownian motion. Furthermore, while the use of a similar distribution is common in the literature on (autoregressive) large language models (e.g., Ziegler et al. (2019)), its application in discrete diffusion models is novel, considering that $p^{\mathrm{pre}}(\cdot)$ cannot be explicitly obtained in our context, unlike autoregressive models.
@@ 4.2 Direct Reward Backpropagation with Gumbel Softmax Trick (DRAKES)
5:           Set $[\pi(t)_{1},\cdots,\pi(t)_{N}]\in\Delta(\mathcal{X})$ where $\pi(t)_{y}=\begin{cases}[\bar{x}^{(i)}_{t-1}]_{y}+\Delta t\sum_{x\in\mathcal{X}}[\bar{x}^{(i)}_{t-1}]_{x}Q^{\theta_{s}}_{y,x}\,(t>0)\\ p_{\text{lim}}(y)\,(t=0)\end{cases}$
@@
$p(x_{t+\Delta t}=y|x_{t}=x)=\mathrm{I}(x=y)+Q^{\theta}_{y,x}(t)(\Delta t).$
Thus, by defining $\pi_{t}=[Q^{\theta}_{1,x}(t)(\Delta t),\cdots,(1+Q^{\theta}_{x,x}(t))\Delta t\cdots,Q^{\theta}_{N,x}(t)(\Delta t)]$, we sample $x_{t+\Delta t}\sim\mathrm{Cat}(\pi_{t})$, where $\mathrm{Cat}(\cdot)$ denotes the categorical distribution.
@@
After approximately sampling from the distribution induced by $P^{\theta_{s}}$, we update the parameter $\theta_{s}$ by maximizing the empirical objective. Although $x_{t}$ itself may not have a valid gradient, $\bar{x}_{t}$ retains the gradient with respect to $\theta$. Therefore, we use the empirical approximation based on $\bar{x}_{t}$. We offer several remarks below, with details in Appendix C:
@@
  • Optimization Objective $g(\theta)$: For the masked diffusion models (1) that we utilized, $g(\theta)$ can be further simplified to reduce computational complexity, as detailed in Appendix C.2.

### v2

@@ 1 Introduction
Diffusion models have gained widespread recognition as effective generative models in continuous spaces, such as image and video generation (Song et al., 2020a; Ho et al., 2022). Inspired by seminal works (e.g., Austin et al. (2021); Campbell et al. (2022); Sun et al. (2022)), recent studies (Lou et al., 2023; Shi et al., 2024b; Sahoo et al., 2024) have shown that diffusion models are also highly effective in discrete spaces, including natural language and biological sequence generation (DNA, RNA, proteins). Unlike autoregressive models commonly used in language modeling, diffusion models are particularly well-suited for biological sequences, where long-range interactions are crucial for the physical behavior of molecules arising from those sequences (e.g., the 3D folded structure of RNA or proteins).
@@
While discrete diffusion models effectively capture conditional distributions (e.g., the distribution of sequences given a specific backbone structure in an inverse protein folding design problem (Dauparas et al., 2022; Campbell et al., 2024)), in many applications, especially for therapeutic discovery, we often aim to generate sequences that are both natural-like and optimize a downstream performance objective. For instance, in the inverse folding problem, we may prefer stable protein sequences (i.e., sequences that fold back into stable protein conformations (Widatalla et al., 2024)); for mRNA vaccine production we desire 5’ UTRs that drive high translational efficiency (Castillo-Hair & Seelig, 2021); for gene and cell therapies, we desire regulatory DNA elements, such as promoters and enhancers, that drive high gene expression only in specific cell types (Taskiran et al., 2024); and for natural language we optimize to minimize harmfulness (Touvron et al., 2023).
@@
To address these challenges, our work introduces a fine-tuning approach for well-pretrained discrete diffusion models that maximizes downstream reward functions. Specifically, we aim to optimize these reward functions while ensuring that the generated sequences maintain a high probability under the original conditional distribution (e.g., the distribution of sequences that fold into a given backbone structure). To achieve this, we formulate the problem as a reward maximization task, analogous to reinforcement learning (RL), where the objective function integrates both the reward terms and the KL divergence with respect to the pre-trained discrete diffusion model, which ensures that the generated sequences remain close to the pre-trained model, preserving their naturalness after fine-tuning. To solve this RL problem, we propose a novel algorithm, DRAKES, that enables direct backpropagation of rewards through entire trajectories by making the originally non-differentiable trajectories differentiable using the Gumbel-Softmax trick (Jang et al., 2016; Maddison et al., 2016).
@@
Our main contribution is an RL-based fine-tuning algorithm, Direct Reward bAcKpropagation with gumbEl Softmax trick (DRAKES), that enables reward-maximizing finetuning for discrete diffusion models (Figure 1). We derive a theoretical guarantee that demonstrates its ability to generate natural and high-reward designs, and demonstrate its performance empirically on DNA and protein design tasks. While similar algorithms exist for continuous spaces (Fan et al., 2023; Black et al., 2023; Uehara et al., 2024b; Venkatraman et al., 2024; Yuan et al., 2023; Guo et al., 2024), our work is the first, to the best of our knowledge, to address these aspects in (continuous-time) discrete diffusion models. This requires addressing unique challenges, as discrete diffusion models are formulated as continuous-time Markov chains (CTMC), which differ from Brownian motion, and the induced trajectories from CTMC are no longer differentiable, unlike in continuous spaces. Our novel theoretical guarantee also establishes a connection with recent advancements in classifier guidance for discrete diffusion models (Nisonoff et al., 2024).
@@ 2 Related Works
Building on the seminal works of Austin et al. (2021); Campbell et al. (2022), recent studies on masked diffusion models (Lou et al., 2023; Shi et al., 2024b; Sahoo et al., 2024) have demonstrated strong performance in natural language generation. Recent advances in masked discrete diffusion models have been successfully applied to biological sequence generation, including DNA and protein sequences (Sarkar et al., 2024; Campbell et al., 2024). Compared to autoregressive models, diffusion models may be particularly well-suited for biological sequences, which typically yield molecules that fold into complex three-dimensional (3D) structures. In contrast to these works, our study focuses on fine-tuning diffusion models to optimize downstream reward functions. One application of our approach is the fine-tuning of protein inverse folding generative models to optimize stability, as discussed in Widatalla et al. (2024). However, unlike this prior work, we employ discrete diffusion models as the generative model.
@@
  • Guidance: Techniques such as classifier guidance (Song et al., 2020b; Dhariwal & Nichol, 2021) and its variants (e.g., Bansal et al. (2023); Chung et al. (2022a); Ho et al. (2022)) introduce gradients from proxy models during inference. However, since gradients are not formally well-defined for discrete states in diffusion, a recent study (Nisonoff et al., 2024) proposed a method specifically designed for discrete diffusion models. Alternative approaches directly applicable to discrete diffusion models include sequential Monte Carlo (SMC)-based methods (Wu et al., 2024; Trippe et al., 2022; Dou & Song, 2024; Cardoso et al., 2023; Phillips et al., 2024). While these guidance-based inference techniques have their own advantages, they generally lead to longer inference times compared to fine-tuned models. We compare our methods against these in terms of generation quality in Section 6.
@@
  • Classifier-free fine-tuning (Ho & Salimans, 2022): This approach constructs conditional generative models, applicable in our setting by conditioning on high reward values. Although not originally designed as a fine-tuning method, it can also be adapted for fine-tuning (Zhang et al., 2023) by adding further controls to optimize. However, in the context of continuous diffusion models, compared to RL-based fine-tuning, several works (Uehara et al., 2024c) have shown that conditioning on high reward values is suboptimal, because such high-reward samples are rare. We will likewise compare this approach to ours in Section 6. Lastly, when pretrained models are conditional diffusion models (i.e., $p(x|c)$) and the offline dataset size consisting of triplets $(c,x,r(x))$ is limited, it is challenging to achieve success. Indeed, for this reason, most current RL-based fine-tuning papers (e.g., Fan et al. (2023); Black et al. (2023); Clark et al. (2023)) do not empirically compare their algorithms with classifier-free guidance.
@@ 3.1 Diffusion Models on Discrete Spaces
In diffusion models, our goal is to model the data distribution $p_{\mathrm{data}}\in\Delta(\mathcal{X})$ using the training data, where $\mathcal{X}$ represents the domain. We focus on the case where $\mathcal{X}=\{1,2,\cdots,N\}$. The fundamental principle is (1) introducing a known forward model that maps the data distribution to a noise distribution, and (2) learning the time reversal that maps the noise distribution back to the data distribution (detailed in Lou et al. (2023); Sahoo et al. (2024); Shi et al. (2024a)).
@@
$\textstyle\frac{dj_{t}}{dt}=Q(t)j_{t},\quad p_{0}\sim p_{\mathrm{data}},$ (1)
where $Q(t)\in\mathbb{R}^{N\times N}$ is the generator. Generally, $j_{t}$ is designed so that $p_{t}$ approaches a simple limiting distribution at $t=T$. A common approach is to add Mask into $\mathcal{X}$ and gradually mask a sequence so that the limiting distribution becomes completely masked (Shi et al., 2024a; Sahoo et al., 2024).
@@
$\textstyle\frac{dj_{T-t}}{dt}=\bar{Q}(T-t)j_{T-t},\quad\bar{Q}_{x,y}(t)=\begin{cases}\textstyle\frac{j_{t}(y)}{j_{t}(x)}Q_{y,x}(t)\,(y\neq x)\\ \textstyle-\sum_{y\neq x}\bar{Q}_{x,y}(t)\,(y=x),\end{cases}$ (2)
where $Q_{x,y}(t)$ is a $(x,y)$-entry of a generator $Q(t)$, representing the transition rate matrix from state $x$ to state $y$. This implies that if we can learn the marginal density ratio $p_{t}(y)/p_{t}(x)$, we can sample from the data distribution at $t=T$ by following the above CTMC controlled by $\bar{Q}(T-t)$. Existing works (e.g., Lou et al. (2023)) demonstrate how to train this ratio from the training data. Especially when we use masked diffusion models (Sahoo et al., 2024; Shi et al., 2024a), we get
$\textstyle\bar{Q}_{x,y}(t)=\begin{cases}\gamma\mathbb{E}[x_{0}=y|x_{t}=\mathrm{Mask}]\quad(y\neq\mathrm{Mask},x_{t}=\mathrm{Mask})\textstyle,\\ -\sum_{z\neq\mathrm{Mask}}\gamma\mathbb{E}[x_{0}=z|x_{t}=\mathrm{Mask}]\quad(y=\mathrm{Mask},x_{t}=\mathrm{Mask})\\ 0\quad(x_{t}\neq\mathrm{Mask})\textstyle\end{cases},$ (3)
@@
When dealing with sequences of length $M$, $x=[x^{\langle 1\rangle},\cdots,x^{\langle M\rangle}]$, we simply consider the factorized rate matrix, i.e., $\bar{Q}_{x,y}=\sum_{i}\bar{Q}_{x,y^{\langle i\rangle}}$ (Campbell et al., 2022), thereby avoiding exponential blowup.
@@ 3.2 Goal: Generating Natural Samples While Optimizing Reward Functions
In our work, we consider a scenario with a pretrained masked discrete diffusion model $p^{\mathrm{pre}}(x|c)\in[\mathcal{C}\to\Delta(\mathcal{X})]$ trained on an extensive dataset and a downstream reward function $r:\mathcal{X}\to\mathbb{R}$. The pretrained diffusion model captures the naturalness or validity of samples. For example, in protein design, $p^{\mathrm{pre}}(\cdot|\cdot)$ could be a protein inverse-folding model that generates amino acid sequences that fold back into the given backbone structure (similar to Campbell et al. (2024)), and $r$ could be a function that evaluates stability. Our objective is to fine-tune a generative model to generate natural-like samples (high $\log p^{\mathrm{pre}}(\cdot|\cdot)$) with desirable properties (high $r(\cdot)$).
@@
We introduce a discrete diffusion model parameterized by $\theta$ from $t=0$ to $t=T$111 Starting from Section 3.2, to simply the notation, we go from $t=0$ to $t=T$ to represent noise to data.:
@@
The parameter $\theta$ from the pretrained model is denoted by $\theta_{\mathrm{pre}}$ and $p_{\text{lim}}$ denotes the initial distribution. The distribution at time $T$ is denoted as $p^{\mathrm{pre}}(\cdot)$, which approximates the training data distribution $p^{\mathrm{data}}$. We denote an element of the generated trajectory from $t=0$ to $t=T$ by $x_{0:T}$. For simplicity, we assume the initial distribution is a Dirac delta distribution (completely masked state), and we often treat the original pretrained diffusion model as an unconditional model for a single token for notational convenience. In this paper, all of the proofs are in Appendix C. We provide an extension to the case when the initial distributions are stochastic in Appendix D.
@@ 4.1 Key Formulation
Perhaps the most obvious starting point for fine-tuning diffusion models to maximize a reward function $r(x_{T})$ is to simply maximize the expected value of the reward under the model’s distribution, i.e., $\mathbb{E}_{x_{0}\sim P^{\theta}}\left[r(x_{T})\right]$, where the expectation is taken over the distribution $P^{\theta}(x_{0:T})$ induced by (4) (i.e., the generator $Q^{\theta}$). However, using only this objective could lead to over-optimization, where the model produces unrealistic or unnatural samples that technically achieve a high reward, but are impossible to generate in reality. Such samples typically exploit flaws in the reward function, for example, by being outside the training distribution of a learned reward or violating the physical assumptions of a hand-engineered physics-based reward (Levine et al., 2020; Clark et al., 2023; Uehara et al., 2024c). We address this challenge by constraining the optimized model to remain close to a pretrained diffusion model, which captures the distribution over natural or realistic samples. More specifically, we introduce a penalization term by incorporating the KL divergence between the fine-tuned model $P^{\theta}(x_{0:T})$ and the pretrained diffusion model $P^{\theta_{\mathrm{pre}}}(x_{0:T})$ in CTMC.
@@
$\exp(r(\cdot)/\alpha)p^{\mathrm{pre}}(\cdot).$ (7)
@@
A similar theorem has been derived for continuous diffusion models (Uehara et al., 2024b, Theorem 1). However, our formulation (5) differs significantly as our framework is based on a CTMC, whereas those works are centered around the Brownian motion. Furthermore, while the use of a similar distribution is common in the literature on (autoregressive) large language models (e.g., Ziegler et al. (2019)), its application in discrete diffusion models is novel, considering that $p^{\mathrm{pre}}(\cdot)$ cannot be explicitly obtained in our context, unlike autoregressive models.
@@ 4.2 Direct Reward Backpropagation with Gumbel Softmax Trick (DRAKES)
5:           Set $[\pi(t)_{1},\cdots,\pi(t)_{N}]\in\Delta(\mathcal{X})$ where $\pi(t)_{y}=\begin{cases}[\bar{x}^{(i)}_{t-1}]_{y}+\Delta t\sum_{x\in\mathcal{X}}[\bar{x}^{(i)}_{t-1}]_{x}Q^{\theta_{s}}_{y,x}\,(t>0)\\ p_{\text{lim}}(y)\,(t=0)\end{cases}\vspace{-6pt}$
@@
$p(x_{t+\Delta t}=y|x_{t}=x)=\mathrm{I}(x=y)+Q^{\theta}_{x,y}(t)(\Delta t).$ (8)
Thus, by defining $\pi_{t}=[Q^{\theta}_{x,1}(t)(\Delta t),\cdots,(1+Q^{\theta}_{x,x}(t))\Delta t\cdots,Q^{\theta}_{x,N}(t)(\Delta t)]$, we sample $x_{t+\Delta t}\sim\mathrm{Cat}(\pi_{t})$, where $\mathrm{Cat}(\cdot)$ denotes the categorical distribution.
@@
After approximately sampling from the distribution induced by $P^{\theta_{s}}$, we update the parameter $\theta_{s}$ by maximizing the empirical objective. Although $x_{t}$ itself may not have a valid gradient, $\bar{x}_{t}$ retains the gradient with respect to $\theta$. Therefore, we use the empirical approximation based on $\bar{x}_{t}$. We offer several remarks below, with details in Appendix E:
@@
  • Optimization Objective $g(\theta)$: For the masked diffusion models (3) that we utilized, $g(\theta)$ can be further simplified to reduce computational complexity, as detailed in Appendix E.2.